# Compress to Impress: Unleashing the Potential of Compressive Memory in Real-World Long-Term Conversations

### v1

@@ Abstract
Existing retrieval-based methods have made significant strides in maintaining long-term conversations. However, these approaches face challenges in memory database management and accurate memory retrieval, hindering their efficacy in dynamic, real-world interactions. This study introduces a novel framework, COmpressive Memory-Enhanced Dialogue sYstems (COMEDY), which eschews traditional retrieval modules and memory databases. Instead, COMEDY adopts a "One-for-All" approach, utilizing a single language model to manage memory generation, compression, and response generation. Central to this framework is the concept of compressive memory, which integrates session-specific summaries, user-bot dynamics, and past events into a concise memory format. To support COMEDY, we curated a large-scale Chinese instruction-tuning dataset, Dolphin, derived from real user-chatbot interactions. Comparative evaluations demonstrate COMEDY’s superiority over traditional retrieval-based methods in producing more nuanced and human-like conversational experiences. Our codes are available at https://github.com/nuochenpku/COMEDY.
Compress to Impress: Unleashing the Potential of Compressive Memory in Real-World Long-Term Conversations
Nuo ${}^{\clubsuit}$  Hongguang ${}^{\diamondsuit}$ Juhua ${}^{\diamondsuit}$  Baoyuan ${}^{\diamondsuit}$  Jia ${}^{\clubsuit}$$^{\clubsuit}$ong Kong University of Science and Technology (Guangzhou) Hong Kong University of Science and Technology ${}^{\diamondsuit}$iaobing.AI nchen022@connect.ust.hk, jialee@ust.hk
@@ 1 Introduction
Maintaining long-term conversations has always been a long-term pursuit in current open-domain dialogue systems Liu et al. (2016); Zhang et al. (2018); Kann et al. (2022); Song et al. (2023), commonly known as chatbots or conversational agents. Long-term conversation refers to the ability of a conversational agent to engage in extended dialogues over multiple interactions, often spanning several days or weeks even months. This setting is challenging because it necessitates not only a deep understanding of the immediate dialogue context but also the retention and integration of key information from past interactions. Effective long-term conversation requires a system to memorize or recall past dialogue snippets, contextual nuances, and user preferences, which are crucial for maintaining coherence and relevance in ongoing interactions Wu et al. (2022); Zhang et al. (2022).
@@
To acquire useful information from past conversations, the most mainstream approach in the field of long-term conversation currently is retrieval-based methods, as illustrated in Figure 1 (a): Firstly, previous works Xu et al. (2022b); Bae et al. (2022) usually employ a memory generator to summarize relevant memories from past sessions, such as user portraits. In this step, the memory generator can either be a separately trained model or a powerful large language model (LLM) like GPT4 OpenAI (2023); Subsequently, a dedicated memory database, or a memory bank, is used to store these memories. Some studies Zhong et al. (2023b) even store past conversational utterances directly in the storage; Going a step further, some works Bae et al. (2022); Wang et al. (2023) propose the use of specific memory management operations to update and iterate the memory database; The final and indispensable step involves employing a sentence-embedding model Guu et al. (2020); Lewis et al. (2020) to retrieve the most relevant memories from the memory database in relation to the current conversation. The current conversation and related memories are then inputted into a specialized response generator to produce the final response.
@@
Despite the notable success achieved by the retrieval-based methods, they encounter several limitations that impact their overall efficacy and applicability: 1) One significant challenge is the unpredictability of performance. The system’s effectiveness is contingent upon several modules (like memory generator and retriever) working in tandem; moreover, the retriever component does not guarantee the retrieval of relevant and effective memories. Sentence-embedding models, commonly used for this purpose, may not always capture the nuances and context of the conversation accurately. 2) Another clear challenge lies in the management of the memory database. As conversations accumulate, the size and complexity of the memory database grow, making it increasingly difficult to manage. Ensuring that the stored information remains relevant and up-to-date is a constant concern, as outdated or irrelevant data can lead to inaccurate or inappropriate responses.
@@
Moreover, current training corpus of long-term conversation chatbots is commonly either involved in constructing personalized dialogue data using LLMs like ChatGPT or hiring crowd-workers to simulate conversations Xu et al. (2022b). Unlike these structured or predictable dialogues, real-world conversations can veer into a wide range of topics, include colloquial language, and incorporate nuanced expressions Chen et al. (2023a). Meanwhile, the memory database in real scenarios needs to store memories from multiple chatbot-users, increasing the difficulty in accurately retrieving relevant memories and maintaining an up-to-date memory database. The above issues present a more pronounced challenge for deploying retrieval-based methods in real-world conversations.
@@
To address these concerns, we propose a LLM-based COmpressive Memory-Enhanced Dialogue sYstems framework (COMEDY). COMEDY marks a significant departure from existing methodologies, as it operates without a retrieval module. At its core, COMEDY adopts a groundbreaking “One-for-All” approach, utilizing a single, unified model to manage the entire process from memory generation, compression to final response generation, as shown in Figure 1 (b): It firstly involves distilling session-specific memory from past dialogues, encompassing fine-grained session summaries, including event recaps, and detailed user and bot portraits; In a break from traditional systems, COMEDY eschews the use of a memory database for storing these insights. Instead, it reprocesses and condenses memories from all past interactions, forming a compressive memory. The first part is the concise events that have occurred throughout all the conversations, creating a historical narrative that the system can draw upon. The second and third parts consist of a detailed user profile and the dynamic relationship changes between the user and chatbot across sessions, both derived from past conversational events. This holistic memory allows COMEDY to generate responses that are not only contextually aware but also personalized and adaptive to the evolving nature of the user-chatbot relationship; Finally, COMEDY skillfully integrates this compressive memory into ongoing conversations, enabling contextually memory-enhanced interactions. Unlike retrieval-based systems that may struggle to fetch pertinent memories from a vast database, COMEDY’s compressive memory is inherently designed to prioritize salient information, allowing for quicker and more accurate memory utilization.
@@
To ensure that COMEDY is well-suited for real-world long-term conversations and overcome the issues of lacking relevant labeled data, we have methodically assembled an large-scale instruction-tuning dataset from actual online user-chatbot interactions, named Dolphin. This dataset contains three tasks: Session-Level Memory Summarization; Memory Compression; Memory-Grounded Response Generation, comprising an extensive collection of 100k samples. Dolphin is well-annotated to support each critical phase in COMEDY’s operation, from memory extraction and compression to integration and response generation. This dataset lays a robust foundation for enhancing COMEDY’s dialogue capabilities, ultimately leading to a more nuanced and human-like conversational experience compared to retrieval-based baselines. Our COMEDY has been deployed on the X Eva platform and has received over 10 million calls, indicating its widespread use and acceptance.
@@
  • We introduce COMEDY, represents a groundbreaking shift from traditional memory retrieval-based dialogue systems. It not relies on any retriever module and memory database, but generates enhanced, memory-informed responses with compressive memory in terms of comprehensive human evaluation.
@@
  • COMEDY could handle the whole long-term conversation interactions via a singular model, achieving a higher degree of result consistency and predictability, reducing computational overhead, eliminates the need for data transfer between multi-models.
  • We further combine COMEDY with Directly Preference Optimization (DPO) Rafailov et al. (2023) alignment strategy, and propose a simple strategy to mine efficient preferred and dispreferred memory-based responses. COMEDY-DPO shows better ability in generating coherent and contextually memory-grounded responses.
@@ 2 Methodology
In this section, we first overview the problem formulation of long-term conversations in COMEDY-style. Then, we introduce three task definitions and detailed data collection in Dolphin. Last, we present the training and DPO details of COMEDY.
@@ 2.1 Problem Formulation
Table 1: Data statistics of each task in Dolphin. In practical, the amount of collected Task 1 data are much larger than Task 2 and 3. To keep the training balance of data distribution, we align the similar volume of data three tasks.
@@ 2.2 Task and Datasets Collection
The source data in Dolphin originates from X Eva [^1] https://xeva-h5.xiaoice.com/Content/Landing, one of the most popular Chinese AI-User social media platform akin to Character.AI. A distinctive feature of Dolphin is that the AI characters within X Eva are defined by the users themselves. This means that each character can have unique personalities, backgrounds, and conversational traits, as determined by the user’s input and creativity.
In the creation of the Dolphin dataset for COMEDY, we first select the episode $D$ that contains 15 sessions at least between same user and AI characters as our source dialogue data, with filtering useless and toxic information. Then we have adopted an efficient LLM-Annotators hybrid approach to annotate each task data Chen et al. (2023d): (1) We initiate the dataset generation using GPT4-Turbo, specifically tailored for dialogue summaries and memory-grounded dialogues. This step is crucial for creating a comprehensive base of dialogues, encompassing a wide range of conversational scenarios and memory contexts; (2) Following the initial generation, three skilled annotators meticulously review and refine the data. This involves correcting inaccuracies, enhancing dialogue quality. The annotators play a vital role in bridging the gap between automated generation and the nuanced understanding required for high-quality COMEDY.
To protect user privacy, all personal identifiers are removed from the dataset. This includes names, locations, or any specific details that could lead to the identification of individuals. Relevant discussion are presented in Ethical Claims.
@@
In the process of gathering data for Task 1, we encounter a substantial challenge. The initial collection yielded over 500,000 session-level data points, making it impractical to annotate all of them through GPT4-Turbo and manual methods due to the sheer volume. To tackle this, we initially focus on annotating a subset of approximately 40,000 data: For each dialogue session in the same episode $D$, we first require the GPT4-Turbo to extract session-level memories, including the event, user and bot portraits in natural sentences. Then annotators edit the generated summaries by add missing information or revising erroneous sentences, resulting session-level memory $m_{n}$. Utilizing the annotated subset, we then develop a specialized LLM for session-level memory generation, efficiently expanding our dataset while maintaining the quality and consistency of the session-level memory annotations across the larger dataset. Samples with none informative content, leading to ineffective memory outputs from LLM or GPT4-Turbo, are filtered out to maintain data quality. As a result, in this task, we collect fine-grained memories $M=\{m_{1},m_{2},\ldots,m_{n}\}$ for each session in $D$.
@@
In this task, the focus is on memory compression, a crucial step in refining the data for COMEDY. GPT4-Turbo is tasked with summarizing all session-level memory $M$ in the episode from Task 1. The output from GPT4-Turbo includes: 1) A Comprehensive User Profile: Detailing characteristics, behavioral patterns, and recent states of the user. 2) Evolving Dynamics between User and Bot: Capturing the relationship’s progression and interaction nuances. 3) Concise Record of Past Events: Summarizing key happenings and dialogues from previous sessions. Considering the potential complexity and variance in the summarization process, GPT4-Turbo is configured to generate outputs three times with a temperature setting of 0.9. This setting allows for a balance between creativity and relevance, enabling GPT4-Turbo to produce diverse and insightful summaries. Then annotators step in to refine and calibrate the outputs, includes: Correcting any inaccuracies or inconsistencies in the summaries; Ensuring that the summarized data accurately reflects the user profiles, relationship dynamics, and event records; Enhancing clarity and conciseness where necessary. This hybrid approach ensures that compressive memory $\hat{M}$ meets the high-quality standards required for the subsequent stages of COMEDY’s development. We show examples of $\hat{M}$ in Table LABEL:table:task2example.
@@
In this task, the process begins with integrating the compressive memory $\hat{M}$, obtained from Task 2, with the incoming conversation at time step t, denoted as $D_{t}$. The annotation process are similar with previous tasks: Initial response drafts are generated by GPT4-Turbo, based on the integrated data of $\hat{M}$ and $D_{t}$. Annotators then review and refine these responses, focusing on aspects like relevance, coherence, and personalization. They ensure that each annotated response $c_{t+1}$ accurately reflects the user’s current state and previous interactions, maintaining high memorability and engagingness. To ensure the scale of the training data, we annotate all sessions within one day closest to the previous $D$ timing as the corpus of Task 3.
Figure 2: The overview training pipeline of COMEDY.
Test Set
@@
Of note, we also manually annotate the session-level memory and the resulting compressive memory in the first 15 sessions. They are used to evaluate the model’s performance in Task 1 and 2. Our Quality Control process in Appendix B, and prompts of each task that takes as inputs into GPT4-Turbo are shown in Appendix E.
As a result, the statistics of our dataset are shown in Table 1. Dolphin comprises a total of 102,882 samples in training and test sets. Tasks 1 and 2 (Memory Extraction and Compression) contain 39,999 and 30,695 samples in training, making up a significant portion of the dataset. Task 3, which involves generating responses based on the compressive memory, comprises 31,131 dialogue sessions. A notable feature of the Dolphin dataset is its inclusion of data from 3,998 different AI characters. The diverse character data ensures that COMEDY is well-equipped to interact with various user personalities and preferences, enhancing its adaptability and realism in user interactions.
@@ 2.3 COMEDY
In practical, we adopt a mixed task training approach to develop COMEDY. This involves simultaneously training the model on the three tasks - session-level memory summarization, memory compression, and memory-grounded response generation - present in the Dolphin dataset. This integration presents the model with a holistic view of the conversation process, from initial memory extraction to final response generation. We utilize the common language modeling objective in SFT, terming the resulting model as $\mathcal{M}(\theta)_{\text{sft}}$.
@@
Compressive Memory-based
@@
Table 3: Human evaluation results of scoring in Task 3: memory-grounded response generation. For COMEDY-GPT4/ChatGPT, the compressive memories are generated by COMEDY-13B.
@@
In order to align the model generating more coherent and contextually appropriate memory-grounded responses, we employ Direct Preference Optimization (DPO) Rafailov et al. (2023) strategy in Task 3. DPO aims to distill a referential SFT policy $\mathcal{M}(\theta)_{\text{sft}}$ by polarizing the preference. Specifically, DPO involves an input labeled pairs ($Y_{w},Y_{l}$) where $Y_{w}$ and $Y_{l}$ denotes the preferred and dispreferred completion. When extended DPO in Memory-grounded generation, the question is: how we obtain the $Y_{w}$ and $Y_{l}$?
To solve this, we propose a simple strategy to automatically construct useful $Y_{w}$ and $Y_{l}$ responses without human annotation. Suppose $\hat{M}$ and $D_{t}$ are given, we ask the GPT4-Turbo to generate the response $Y_{w}$ must align the $\hat{M}$. Meanwhile, we also require GPT4-Turbo to generate the response $Y_{l}$ that is totally against the $\hat{M}$. For example, the prompts are illustrated like: “If $\hat{M}$ shows users like something, you should generate the response with the meaning of users hates it…”. Thus, the overall training objective of DPO can be formalized as:
@@
where x is the concatenation of $\hat{M}$ and $D_{t}$, $\beta$ is a hyper-parameter. The overview of our training pipeline is shown in Figure 2 and training instruction in Appendix E.

### v2

# Beyond Retrieval: Embracing Compressive Memory in Real-World Long-Term Conversations

@@ Abstract
Existing retrieval-based methods have made significant strides in maintaining long-term conversations. However, these approaches face challenges in memory database management and accurate memory retrieval, hindering their efficacy in dynamic, real-world interactions. This study introduces a novel framework, COmpressive Memory-Enhanced Dialogue sYstems (COMEDY), which eschews traditional retrieval modules and memory databases. Instead, COMEDY adopts a “One-for-All” approach, utilizing a single language model to manage memory generation, compression, and response generation. Central to this framework is the concept of compressive memory, which integrates session-specific summaries, user-bot dynamics, and past events into a concise memory format. To support COMEDY, we collect the biggest Chinese long-term conversation dataset, Dolphin, derived from real user-chatbot interactions. Comparative evaluations demonstrate COMEDY’s superiority over traditional retrieval-based methods in producing more nuanced and human-like conversational experiences.
Beyond Retrieval: Embracing Compressive Memory in Real-World Long-Term Conversations
Nuo Chen  Hongguang Li Juhua Huang  Baoyuan Wang  Jia Li Hong Kong University of Science and Technology (Guangzhou) Hong Kong University of Science and Technology Xiaobing.AI nchen022@connect.ust.hk, jialee@ust.hk
@@ 1 Introduction
Maintaining long-term conversations has always been a long-standing pursuit in current open-domain dialogue systems Liu et al. (2016); Zhang et al. (2018); Kann et al. (2022), commonly known as chatbots or conversational agents. Long-term conversation refers to the ability of a conversational agent to engage in extended dialogues over multiple interactions, often spanning several days, weeks, or even months. This setting is challenging because it necessitates not only a deep understanding of the immediate dialogue context but also the retention and integration of key information from past interactions. Effective long-term conversation requires a system to memorize or recall past dialogue snippets, contextual nuances, and user preferences, which are crucial for maintaining coherence and relevance in ongoing interactions Wu et al. (2022); Zhang et al. (2022).
@@
To acquire useful information from past conversations, the most mainstream approach in the field of long-term conversation currently is retrieval-based methods, as illustrated in Figure 1 (a): Firstly, previous works Xu et al. (2022b); Bae et al. (2022) usually employ a memory generator to summarize relevant memories from past sessions, such as user portraits; Subsequently, a dedicated memory database, or a memory bank, is used to store these memories. Some studies Zhong et al. (2023b) even store past conversational utterances directly in the storage; Going a step further, some works Bae et al. (2022); Wang et al. (2023b) propose the use of specific memory management operations to update and iterate the memory database; The final and indispensable step involves employing a sentence-embedding model Guu et al. (2020); Lewis et al. (2020) to retrieve the most relevant memories from the memory database in relation to the current conversation. The current conversation and related memories are then inputted into a specialized response generator to produce the final response.
@@
Despite the notable success achieved by retrieval-based methods, they encounter several limitations that impact their overall efficacy and applicability: 1) One significant challenge is the unpredictability of the performance. The system’s effectiveness is contingent upon several modules (like memory generator and retriever) working in tandem; moreover, the retriever component does not guarantee the retrieval of relevant and effective memories. Sentence-embedding models Gao et al. (2021); Reimers and Gurevych (2019), commonly used for this purpose, may not always capture the nuances and context of the conversation accurately. 2) Another clear challenge lies in the management of the memory database. As conversations accumulate, the size and complexity of the memory database grow, making it increasingly difficult to manage. Ensuring that the stored information remains relevant and up-to-date is a constant concern, as outdated or irrelevant data can lead to inaccurate or inappropriate responses.
@@
Moreover, current training corpus of long-term conversation chatbots is commonly either involved in constructing personalized dialogue data using LLMs Wang et al. (2023a) like ChatGPT or hiring crowd-workers to simulate conversations Xu et al. (2022b). Unlike these structured or predictable dialogues, real-world conversations can veer into a wide range of topics, include colloquial language, and incorporate nuanced expressions Chen et al. (2023a). Ensuring that a retrieval model is robust enough to handle such real-world variations in language can be extremely difficult. Meanwhile, the memory database in real scenarios needs to store memories from multiple chatbot-users, increasing the difficulty in accurately retrieving relevant memories and maintaining an up-to-date memory database. The above issues present a more pronounced challenge for deploying retrieval-based methods in real-world conversations.
@@
To address these concerns, we propose a LLM-based COmpressive Memory-Enhanced Dialogue sYstem (COMEDY). COMEDY marks a significant departure from existing methodologies, as it operates without a retrieval module. At its core, COMEDY adopts a groundbreaking “One-for-All” approach, utilizing a single, unified model to manage the entire process from memory generation, compression to final response generation, as shown in Figure 1 (b): It firstly involves distilling session-specific memory from past dialogues, encompassing fine-grained session summaries, including event recaps, and detailed user and bot portraits; In a break from traditional systems, COMEDY eschews the use of a memory database for storing these insights. Instead, it reprocesses and condenses memories from all past interactions, forming a compressive memory, including concise events, a detailed user profile and dynamic relationship changes between the user and chatbot across past sessions. This holistic memory allows COMEDY to generate responses that are not only contextually aware but also personalized and adaptive to the evolving nature of the user-chatbot relationship; Finally, COMEDY skillfully integrates this compressive memory into ongoing conversations, enabling contextually memory-enhanced interactions. Unlike retrieval-based systems that may struggle to fetch pertinent memories from a vast database, COMEDY’s compressive memory is inherently designed to prioritize salient information, allowing for quicker and more accurate memory utilization.
@@
To ensure that COMEDY is well-suited for real-world long-term conversations and overcome the issues of lacking relevant labeled data, we have methodically assembled a large-scale instruction-tuning dataset from actual online user-chatbot interactions, named Dolphin. This dataset contains three tasks: Session-Level Memory Summarization; Memory Compression; Memory-Grounded Response Generation, comprising an extensive collection of 100k samples. Dolphin is well-annotated to support each critical phase in COMEDY’s operation, from memory extraction and compression to integration and response generation. This dataset lays a robust foundation for enhancing COMEDY’s dialogue capabilities, ultimately leading to a more nuanced and human-like conversational experience compared to retrieval-based baselines.
@@
  • We introduce a new framework, named COMEDY, represents a groundbreaking shift from traditional memory retrieval-based dialogue systems. It does not rely on any retriever module or memory database, but generates enhanced, memory-informed responses with compressive memory.
@@
  • COMEDY could handle the whole long-term conversation interactions via a singular model, achieving a higher degree of result consistency and predictability, reducing computational overhead, and eliminating the need for data transfer between multi-models.
@@ 2 Methodology
In this section, we first overview the problem formulation of long-term conversations in COMEDY-style. Then, we introduce three task definitions and detailed data collection in Dolphin. Last, we present the training strategies of COMEDY.
@@ 2.1 Problem Formulation
Table 1: Data statistics for each task in Dolphin. In practice, the amount of collected Task 1 data is much larger than Task 2 and 3. To keep the training balance of data distribution, we align the similar volume of data in three tasks.
@@ 2.2 Task and Datasets Collection
The source data in Dolphin originates from X Eva [^1] https://xeva-h5.xiaoice.com/Content/Landing, one of the most popular Chinese AI-User social media platforms akin to Character.AI. A distinctive feature of Dolphin is that the AI characters within X Eva are defined by the users themselves. This means that each character can have unique personalities, backgrounds, and conversational traits, as determined by the user’s input and creativity.
In the creation of the Dolphin dataset for COMEDY, we first select the episode $D$ that contains at least 15 sessions between the same user and AI characters as our source dialogue data after filtering out useless and toxic information. Then we adopt an efficient LLM-Human Annotators hybrid approach to annotate each task data Chen et al. (2023d): (1) We initiate the dataset annotation using GPT4-Turbo, specifically tailored for dialogue summaries and memory-grounded dialogues. This step is crucial for creating a comprehensive base of dialogues, encompassing a wide range of conversational scenarios and memory contexts; (2) Following the initial generation, three skilled annotators meticulously review and refine the data. This involves correcting inaccuracies, enhancing dialogue quality. The annotators play a vital role in bridging the gap between automated generation and the nuanced understanding required for high-quality COMEDY.
To protect user privacy, all personal identifiers are removed from the dataset. This includes names, locations, or any specific details that could lead to the identification of individuals. Relevant details are presented in Appendix B.
@@
In the process of gathering data for Task 1, we encounter a substantial challenge. The initial collection yielded over 500,000 session-level data points, making it impractical to annotate all of them through GPT4-Turbo and manual methods due to the sheer volume. To tackle this, we initially focus on annotating a subset of approximately 40,000 data: For each dialogue session in the same episode $D$, we first require the GPT4-Turbo to extract session-level memories, including the event, user and bot portraits in natural sentences. Then annotators edit the generated summaries by adding missing information or revising erroneous sentences, resulting in session-level memory $m_{n}$. Utilizing the annotated subset, we then develop a specialized LLM for session-level memory generation, efficiently expanding our dataset while maintaining the quality and consistency of the session-level memory annotations across the larger dataset. Samples with no informative content, leading to ineffective memory outputs from LLM or GPT4-Turbo, are filtered out to maintain data quality. As a result, in this task, we collect fine-grained memories $M=\{m_{1},m_{2},\ldots,m_{n}\}$ for each session in $D$.
@@
In this task, the focus is on memory compression. GPT4-Turbo is tasked with summarizing all session-level memory $M$ in the episode from Task 1, outputting the compressive memory $\hat{M}$. It includes: 1) A Comprehensive User Profile: Detailing characteristics, behavioral patterns, and recent states of the user. 2) Evolving Dynamics between User and Bot: Capturing the relationship’s progression and interaction nuances. 3) Concise Record of Past Events: Summarizing key happenings and dialogues from previous sessions. Considering the potential complexity and variance in the summarization process, GPT4-Turbo is configured to generate outputs three times with a temperature setting of 0.9. This setting allows for a balance between creativity and relevance, enabling GPT4-Turbo to produce diverse and insightful summaries. Then annotators step in to refine and calibrate the outputs, which includes: Correcting any inaccuracies or inconsistencies in the summaries; Ensuring that the summarized data accurately reflects the user profiles, relationship dynamics, and event records; Enhancing clarity and conciseness where necessary. This hybrid approach ensures that compressive memory $\hat{M}$ meets the high-quality standards required for the subsequent stages of COMEDY’s development. We show examples of $\hat{M}$ in Table 10.
@@
Similarly, given compressive memory $\hat{M}$ and incoming conversation $D_{t}$, GPT4-Turbo outputs the memory-based responses. Annotators then review and refine these responses, focusing on aspects like relevance, coherence, and personalization. They ensure that each annotated response $c_{t+1}$ accurately reflects the user’s current state and previous interactions, maintaining high memorability and engagingness. To ensure the scale of the training data, we annotate all sessions within one day closest to the previous $D$ timing as the corpus of Task 3.
Test Set.
@@
Quality Control.
Ensuring high-quality data is paramount for the accuracy, reliability, and overall performance of the system. In this work, we employ several strategies to control data quality:
  • Annotator Performance Monitoring: Regular assessments of annotator performance are conducted to ensure consistent quality across the team (every day). This includes evaluating their accuracy, attention to detail, and adherence to annotation guidelines.
Note, we also manually annotate the session-level memory and the resulting compressive memory in the first 15 sessions. They are used to evaluate the model’s performance in Task 1 and 2. Our prompts and examples of each task are shown in Appendix E, Table 7-10.
Statistics.
As a result, the statistics of our dataset are shown in Table 1. Dolphin comprises a total of 102,882 samples in training and test sets. Tasks 1 and 2 (Memory Extraction and Compression) contain 39,999 and 30,695 samples in training, respectively, making up a significant portion of the dataset. Task 3, which involves generating responses based on the compressive memory, comprises 31,131 dialogue sessions. A notable feature of the Dolphin dataset is its inclusion of data from 3,998 different AI characters. The diverse character data ensures that COMEDY is well-equipped to interact with various user personalities and preferences, enhancing its adaptability and realism in user interactions.
@@ 2.3 COMEDY
In practice, we adopt a mixed-task training approach to develop COMEDY. This involves simultaneously training the model on the three tasks - session-level memory summarization, memory compression, and memory-grounded response generation - present in the Dolphin dataset. This integration presents the model with a holistic view of the conversation process, from initial memory extraction to final response generation. We utilize the common language modeling objective in SFT, terming the resulting model as $\mathcal{M}(\theta)_{\text{sft}}$.
@@
Memory-related
MemoryBank-ChatGPT 1.25 0.94 0.42 0.45 1.52 0.92
Resum-ChatGPT 1.31 0.97 0.47 0.44 1.49 0.93
@@
Table 3: Human scoring evaluation in Task 3: memory-grounded response generation. For COMEDY-GPT4/ChatGPT, the compressive memories are generated by COMEDY-13B.
@@
Experimentally, we find that the SFT model may struggle with maintaining consistency and coherence in generated memories. In order to align the model generating more contextually appropriate memory-grounded responses, we employ Direct Preference Optimization (DPO) Rafailov et al. (2023) strategy in Task 3. DPO aims to distill a referential SFT policy $\mathcal{M}(\theta)_{\text{sft}}$ by polarizing the preference. By polarizing the preferred responses (aligned with memory) and dispreferred responses (against memory), DPO ensures that the generated outputs remain consistent with the user’s past interactions and the overall context of the conversation. Specifically, DPO involves input labeled pairs ($Y_{w},Y_{l}$) where $Y_{w}$ and $Y_{l}$ denotes the preferred and dispreferred completion. When extended DPO in Memory-grounded generation, the question is: how we obtain the $Y_{w}$ and $Y_{l}$?
To solve this, we propose a simple strategy to automatically construct useful $Y_{w}$ and $Y_{l}$ responses without human annotation. Suppose $\hat{M}$ and $D_{t}$ are given, we ask the GPT4-Turbo to generate the response $Y_{w}$ must align the $\hat{M}$. Meanwhile, we also require GPT4-Turbo to generate the response $Y_{l}$ that is totally against the $\hat{M}$. For example, the prompts are illustrated like: “If $\hat{M}$ shows users like something, you should generate the response with the meaning of users hate it…”, shown in Table 6. Formally, the training objective of DPO is:
@@
where x is the concatenation of $\hat{M}$ and $D_{t}$, $\beta$ is a hyperparameter. Training instructions are in Appendix E.